Hello
I was wondering if you would be able to assist me please. I am given a quadratic problem, and was hoping you could show me how to set this out from the start. The worded ones always trouble me.
It goes like this:
If the profit $P for manufacturing x articles is given by P=140+60x-3x^2. How many such articles (x) should be manufactured for maximum profit? How much is the profit?
Thank you so very much.

Answer by scott8148(6628)

the maximum value for P lies on the axis of symmetry of the graph of the function

for an equation of the form ax^2+bx+c, the equation of the axis of symmetry is x=-b/2a

x=-6/(2(-3)) __ x=10 __ this is the number of articles for maximim profit
__ substitute this value into the original equation to find P